Numerical Example

A representative firm is operating in a perfectly competitive industry. The firm’s total cost, TC, is given by the equation TC = 50 + 5q2 , where q is output. Based on this equation, the marginal cost, MC, is 10q.

1. If the output price is $100, what is the short-run profit-maximizing output?

2. How much profit does this firm make at that level of output?

3. What do you expect to happen in the market in the long run?
Answer #1

Answer

1)

The firm produces at MC=P

10q=100

q=10 units

the firm produces 10 units

=======

2)

total revenue =P*q=10*100=1000

TC=50+5*10^2

=550

Profit=TR-TC=1000-550=$450

the profit is $450

3)

the firm make a profit so the firm attracts new firms in the market in the long run-up to the profit is zero.